OpenStudy (anonymous):

Find theta if cosecant of theta =2.264

OpenStudy (anonymous):

\[\csc(\theta)=\frac{ 1 }{ \sin(\theta) }\]From that concept, change \[\csc(\theta)\]Then, you get \[\frac{ 1 }{ \sin (\theta) }=2.264\]\[\sin(\theta)=\frac{ 1 }{ 2.264 }\]

OpenStudy (anonymous):

Thank you